### v2

# Nonlinear Online Optimization for Vehicle-Home-Grid Integration including Household Load Prediction and Battery Degradation

Abstract
This paper investigates the economic impact of vehicle-home-grid integration, by proposing an online energy management algorithm that optimizes energy flows between an electric vehicle (EV), a household, and the electrical grid. The algorithm leverages vehicle-to-home (V2H) for self-consumption and vehicle-to-grid (V2G) for energy trading, adapting to real-time conditions through a hybrid long short-term memory (LSTM) neural network for accurate household load prediction, alongside a comprehensive nonlinear battery degradation model accounting for both cycle and calendar aging. Simulation results reveal significant economic advantages: compared to smart unidirectional charging, the proposed method yields an annual economic benefit of up to €3046.81, despite a modest 1.96% increase in battery degradation. Even under unfavorable market conditions, where V2G energy selling generates no revenue, V2H alone ensures yearly savings of €425.48. A systematic sensitivity analysis investigates how variations in battery capacity, household load, and price ratios affect economic outcomes, confirming the consistent benefits of bidirectional energy exchange. These findings highlight the potential of EVs as active energy nodes, enabling sustainable energy management and cost-effective battery usage in real-world conditions.
I INTRODUCTION
In recent years, the widespread adoption of electric vehicles (EVs) has experienced a significant growth, bringing major changes to the transportation sector. EVs are providing a more sustainable solution to the challenges posed by climate change. However, this transition is also having a significant impact on global energy demand. With the exponential growth in the number of EVs on the road, the need for energy to recharge their batteries increases daily, placing significant challenges on electric grids. Projections indicate that global electricity demand for EV charging is expected to increase by 633% from 2023 to 2030 [1].
The primary goal of traditional charging systems is to charge the battery once an EV is connected to the power socket [2]. However, a vehicle is parked approximately 96% of the time [3]. This highlights the potential to leverage EVs for active interaction with electric grids. With the emergence of bidirectional power transfer for EVs, it is now possible to reduce the costs for EV owners while providing a service to the grid. In this context, concepts such as vehicle-to-grid (V2G) and vehicle-to-home (V2H) are emerging as groundbreaking solutions, where EVs are no longer merely means of transportation but actual energy nodes. V2G allows EVs to supply energy back to the grid, providing services such as load balancing and frequency regulation [4], while allowing the EV owner to generate profit by selling stored energy. V2H allows EVs to supply power to a home, supporting home energy management and enabling the user to reduce energy costs and increase self-sufficiency [5]. In addition to being used to minimize electricity consumption costs, the V2H can also be used as a backup source for the load connected to the home in the event of a power failure from the grid [4].
The adoption of V2G and V2H technologies largely depends on user acceptance and preferences. A survey conducted in Sweden highlights that vehicle-home-grid integration presents both challenges and opportunities [6]. First, users tend to prefer V2H systems, perceiving them as relatively straightforward. This preference is primarily driven by the immediate financial benefits (i.e., energy cost savings for household consumption) and the added resilience provided during grid outages. While V2G is also seen as promising for its various services to the grid and the offered opportunity for energy arbitrage, this technology presents great challenges, particularly regarding battery longevity and the risk of insufficient charge capacity for daily vehicle use.
The user preferences highlighted in [6] underline the importance of addressing both technical and economic challenges to enhance the adoption of bidirectional charging systems. To this end, considerable research effort has been devoted to developing energy management algorithms for V2G and V2H.
In the relevant literature, most approaches rely on offline optimization. For example, a multi-integer linear programming (MILP) framework was developed to optimize the operation of smart households using V2G and V2H, aiming to minimize user costs [7]. Without considering V2H, Dong et al. formulated the day-ahead V2G scheduling problem within the multi-agent reinforcement learning framework to optimize the peak shaving performance for the electric grid [8]. V2G and V2H applications for energy trading were also explored in [9], which reports that an average German household with a photovoltaic system, heat pump, and stationary battery can generate annual revenues of approximately €310. However, these three works ignored battery degradation in the optimization for simplicity, which tend to result in suboptimal solutions due to excessive battery costs. To mitigate the risk, Khezri et al. presented a MILP-based optimal scheduling model for EVs, in which V2G and a linearized battery degradation model were considered [10]. Following this trend, Lee et al. [11] presented a linear programming-based V2G optimization focusing on frequency regulation according to the rules of the Swedish market. A common feature of these works is that they assume full knowledge of input data in advance, such as household loads, and then solve the problem offline, making them less applicable to real-world with dynamic scenarios.
To mitigate the potential drawbacks of offline optimization, recent research has increasingly focused on optimizing V2G technology online. Li et al. [12] proposed an online battery anti-aging V2G scheduling method and utilized fuzzy logic control to solve the optimization in two stages. Therein, only the cycle aging was considered during offline calibration to set the fuzzy control parameters, whereas battery degradation attributed to both cycle aging and calendar aging was not explicitly considered during the online phase. In [13], particle swarm optimization (PSO) was used to solve the V2G optimization with a sliding window; as for the battery degradation, it only considers the cycle aging with the rain-flow cycle counting method. Although battery degradation has been considered, the existing V2G optimization algorithms heavily rely overly simplified degradation models that fail to capture calendar and cycle aging comprehensively. Furthermore, these online algorithms have not systematically synergized V2G, V2H, battery dynamics, and the inherent uncertainties associated with household energy consumption. These limitations highlight the need for more practical and accurate solutions that can adapt to real-time conditions while mitigating the impact on battery lifespan.
To address the identified research gaps, this paper proposes a nonlinear online optimization algorithm for the vehicle-home-grid integration, aiming to minimize user costs through energy trading based on hourly price variations. This algorithm is carried out by considering a single user, who owns both an EV and a house. To obtain more predictable results in real-world scenarios, the algorithm is implemented in an online manner, meaning that data is processed as it arrives rather than being known in advance. Specifically, we introduce a hybrid long short-term memory (LSTM) neural network to predict the domestic load, enhancing real-time energy allocation. Additionally, we use a detailed battery model to account for both calendar and cycle aging, particularly the nonlinear aging characteristics in response to various coupled stress factors. As a result, the proposed algorithm minimizes the overall user electricity and battery costs, contributing to more sustainable and economically viable energy management.
II METHODOLOGY
The considered vehicle-home-grid integration, as illustrated in Fig. 1, has three actors: an EV, a house, and the power grid. The EV can supply energy to the grid and the house (V2G and V2H), while grid-to-vehicle (G2V) represents the energy used to recharge the EV. The house, on the other hand, can also receive energy in a standard way, directly from the power grid through grid-to-home (G2H).
Figure 1: Vehicle-home-grid interaction of the proposed algorithm. The arrows indicate the allowable energy flows for V2H, V2G, G2V and G2H.
As most EV users prefer to charge their EV at home rather than at public charging points [14], the present work assumes that the EV is only charged when parked at home.
Since the model is implemented online, the exact time when the user parks the EV is not known in advance. However, once the EV is parked, the user communicates the time at which they will pick it up again.
The electricity price is known in advance, thanks to the day-ahead pricing. However, the household load is only known up to the current time, while future values remain uncertain. To address this uncertainty, an online hybrid LSTM neural network is used to predict the future household load. Additionally, at each time step a correction algorithm refines the predicted values based on actual consumption, making it possible for the control algorithm to calculate the optimal energy flows within the given time window.
II-A Vehicle-Home-Grid Control
While the EV is parked, an optimization problem is formulated to minimize the user cost for energy trading and battery degradation:
$\min\sum_{t}EC_{t}+BC_{t}+s_{t},$ (1)
where $t$ represents time, $s_{t}$ is a slack variable for a soft constraint related to the EV’s SoC at the pickup time, and $EC_{t}$ and $BC_{t}$ denote the energy and battery costs, respectively.
$EC_{t}$ and $BC_{t}$ can be calculated by
$EC_{t}=\>$ $(E^{G2V}_{t}+E^{G2H}_{t})\cdot p_{t}-E^{V2G}_{t}\cdot\gamma\cdot p_{t},$ (2)
$BC_{t}=\>$ $NV\cdot\frac{BD_{t}(\%)}{100\%-EoL(\%)},$ (3)
where the energy flows are expressed in kWh, and $p_{t}$ is the day-ahead energy price, expressed in €/kWh. Energy costs arise from purchasing energy from the grid for G2V and G2H, minus the profits from selling energy to the grid (V2G). $\gamma$ is the price ratio, representing the ratio between the selling price and the buying price of energy. If $\gamma=1$, it means the buying and selling prices are equal. In (3), $BD_{t}$ is the battery degradation in percentage of the initial battery capacity, $EoL$ is the battery capacity at the end of life, which is assumed to be 80%, and $NV$ is the net value of the battery.
Based on economic principles discussed in [15], $NV$ is calculated through
$NV=C_{rep}\cdot\frac{1}{(1+i_{r})^{L}}-C_{rv}\cdot\frac{1}{(1+i_{r})^{L}},$ (4)
where $C_{rep}$ is the battery replacement cost, $C_{rv}$ is the battery residual value, $i_{r}$ is the yearly discount rate, and $L$ represents the nominal battery life in years.
The objective function in (1) is subject to the following set of constraints:
$E^{V2G}_{t},E^{V2H}_{t},E^{G2V}_{t},E^{G2H}_{t},s_{t}\geq\>$ $0,$ (5)
$E^{G2V}_{t}\leq\>$ $E_{\max},$ (6)
$E^{V2G}_{t}+E^{V2H}_{t}\leq\>$ $E_{\max},$ (7)
$0\%\leq SoC_{t}\leq\>$ $100\%,$ (8)
$SoC_{t}=SoC_{t-1}+\frac{E^{G2V}_{t}}{E_{b}}-\>$ $\frac{E^{V2G}_{t}}{E_{b}}$
$-\frac{E^{V2H}_{t}}{E_{b}}-\frac{D}{D_{r}},\>\>\>\>$ (9)
$SoC_{t}+s_{t}\geq\>$ $SoC^{goal}_{t},$ (10)
$HL_{t}=\>$ $E^{G2H}_{t}+E^{V2H}_{t},$ (11)
$E^{G2V}_{t}+E^{G2H}_{t}\leq\>$ $G_{t}.$ (12)
In (5), all energy flows and $s_{t}$ must be non-negative. The input energy $E^{G2V}_{t}$ and the output energy from the EV ($E^{V2G}_{t}+E^{V2H}_{t}$) cannot exceed the maximum limit $E_{\max}$, imposed by the EV charger, resulting in the constraints formulated in (6) and (7). The constraint (8) ensures that the EV’s SoC ranges from 0 to 100%. The SoC evolution in (9) depends on the previous SoC, the energy exchanged with the battery (normalized by its capacity, $E_{b}$), and the distance traveled (normalized by the battery range, $D_{r}$). Notice that, while the EV is parked, $D=0$. (10) is a soft constraint ensuring the EV reaches the desired SoC of the user. Specifically, it requires the SoC to be at least $SoC^{goal}_{t}$, which is 80% at pickup times and zero otherwise. If the constraint is not met, $s_{t}$ compensates for the deviation, contributing to an increase in the objective function (1). The household load $HL_{t}$ is met by energy supplied either from the grid or the EV, according to (11). However, the total energy purchased from the grid cannot exceed the available supply $G_{t}$ (which is here assumed to be sufficiently large to always ensure the constraint (12) is never active).
Note that (1)–(12) apply only when the EV is parked. However, the model simulates a realistic scenario where the EV alternates between parking and driving throughout the simulation. During driving, no optimization is needed, as the EV is absent and all its energy flows are zero. In this phase, the SoC depends on the previous step and the distance traveled $D$, while the household load is fully supplied by the grid ($E_{t}^{G2H}$).
II-B Battery Model
An empirical battery model is employed to describe the degradation characteristics, including both calendar and cycle aging. This model has been experimentally validated against real-world lithium-iron-phosphate batteries [16] and has been widely applied in the literature [17]. While the details of the model can be found in [16], we present the key equations here for calculating the battery degradation in the vehicle-home-grid control.
The calendar aging $BD_{t}^{cal}$ is a function of temperature $T$, SoC and time in hours, as computed by
$BD_{t}^{cal}=\frac{K^{cal}_{t}(T,SoC)}{2\sqrt{t}}\Delta t,$ (13)
with $t$ being the cumulative time and $\Delta t$ the sample time, assumed to be one hour in this work. $K^{cal}_{t}(\cdot,\cdot)$ is a stress factor which depends on $T$ and SoC according to
$K^{cal}_{t}(T$ $,SoC)=k_{cal,ref}\!\cdot\!\exp\left[\frac{-E_{a}^{cal}}{R_{g}}\!\left(\!\frac{1}{T}\!-\!\frac{1}{T_{ref}}\right)\!\right]$
$\cdot\!\left[\!\exp\!\left(\frac{\alpha F}{R_{g}}\frac{U_{a,ref}\!-\!U_{a}(SoC_{t})}{T_{ref}}\!\right)\!+\!k_{0}\!\right],$ (14)
where $U_{a}$ is the anode open-circuit potential (computed as a function of SoC) [16]. We assume a constant $T$ over time, equal to 15°C.
The cycle aging depends on three different conditions: high temperature, low temperature, and low temperature with high SoC. The cycle aging due to high temperature $BD_{t}^{cyc,hT}$ is computed as follows:
$BD_{t}^{cyc,hT}=\frac{K^{cyc,hT}_{t}(T)}{2\sqrt{Q_{tot}}}\Delta Q_{tot},$ (15)
where $Q_{tot}$ is the cumulative total energy throughput (Ah) for charging and discharging, while $\Delta Q_{tot}$ is the instantaneous total energy throughput (in $\Delta t$). $Q_{tot}$ is obtained by summing the energy in input/output (Wh) over time and dividing by the voltage. The stress factor $K^{cyc,hT}_{t}(\cdot)$ is a function of $T$ in the form
$K^{cyc,hT}_{t}\!=k_{cyc,hT,ref}\!\cdot\!\exp\!\left[\!\frac{-E_{a,hT}^{cyc}}{R_{g}}\!\left(\!\frac{1}{T}\!-\!\frac{1}{T_{ref}}\!\right)\!\right].$ (16)
The cycle aging due to low temperature $BD_{t}^{cyc,lT}$ is computed by
$BD_{t}^{cyc,lT}=\frac{K^{cyc,lT}_{t}(T,I_{Ch})}{2\sqrt{Q_{ch}}}\Delta Q_{ch}.$ (17)
Now, $Q_{ch}$ is the cumulative energy throughput (Ah) for charging. The stress factor $K^{cyc,lT}_{t}(\cdot,\cdot)$ is a function of $T$ and the charging current rate $I_{Ch}$ (obtained as $\Delta Q_{Ch}/\Delta t$), which is formulated as
$K^{cyc,lT}_{t}=\>$ $k_{cyc,lT,ref}\exp\left[\frac{-E_{a,lT}^{cyc}}{R_{g}}\left(\frac{1}{T}-\frac{1}{T_{ref}}\right)\right]$
$\cdot\exp\left(\beta_{lT}\frac{I_{Ch}-I_{Ch,ref}}{C_{0}}\right).$ (18)
The cycle aging due to low temperature and high SoC $BD_{t}^{cyc,lThSoC}$ is computed as
$BD_{t}^{cyc,lThSoC}=K^{cyc,lThSoC}_{t}(T,I_{Ch},SoC)\cdot\Delta Q_{ch},$ (19)
where the stress factor $K^{cyc,lThSoC}_{t}(\cdot,\cdot,\cdot)$ is a function of $T$, current rate, and SoC, given by
$\begin{split}&K^{cyc,lThSoC}_{t}\!\!=\!k_{cyc,lThSoC,ref}\exp\!\bigg{[}\!\frac{\!-E_{a,lThSoC}^{cyc}}{R_{g}}\!\left(\!\frac{1}{T}\!-\!\frac{1}{T_{ref}}\!\right)\!\bigg{]}\!\\ &\!\cdot\!\exp\!\left(\!\beta_{lThSoC}\frac{\!I_{Ch}\!-\!I_{Ch,ref}}{C_{0}}\!\right)\frac{\operatorname{sgn}\left(SoC_{t}\!-\!SoC_{ref}\right)\!+\!1}{2}.\end{split}$ (20)
Finally, the total battery degradation $BD_{t}$ in (3), defined as the capacity loss in percentage, is computed at each time step by
$BD_{t}\!=\!BD_{t}^{cal}\!+\!BD_{t}^{cyc,hT}\!+\!BD_{t}^{cyc,lT}\!+\!BD_{t}^{cyc,lThSoC}.$ (21)
In this work, we assume that battery cells in the EV are well-managed, resulting in homogeneous characteristics across different cells. In accordance with a Polestar EV, we set the battery system’s capacity $E_{b}$ to 82 kWh (corresponding to $C_{0}$ = 205 Ah), the nominal voltage to 400 V, and the vehicle’s driving range $D_{r}$ as 514 km. Also, the maximum energy for charging/discharging ($E_{max}$) is set to 11 kWh.
Using the LiFePO4 battery model, the cell open-circuit voltage from [16] is scaled to a nominal voltage of 400 V, as shown in Fig. 2. Since the control algorithm operates in discrete time with a sampling time of one hour, the voltage transients are not considered. Instead, at each time step $t$, we approximate the voltage value by the average between $V(SoC_{t-1})$ and $V(SoC_{t})$.
Figure 2: Voltage profile with respect to SoC for the battery pack.
II-C Household Load Prediction and Management
In real-world applications, the household load is largely affected by weather and user behavior. Due to substantial uncertainties and stochasticity, the load values are typically unknown in advance. This poses challenges for the optimal energy flow to be computed when the EV is parked. To address this, we develop a real-time household load predictor based on historical data and patterns. Specifically, a hybrid long short-term memory neural network (LSTM) is applied. LSTM is very powerful in capturing dependencies in sequential data and is particularly suitable for time series energy consumption data that exhibit recurring daily and seasonal patterns. The hybrid LSTM approach is adopted from [18, 19] to incorporate additional features that influence energy consumption, providing the model with more contextual information and improved prediction accuracy.
The goal of the hybrid LSTM is to predict the household load one hour ahead ($t+1$). From a chosen dataset, a total of four features were extracted:
  1. $f1)$
    energy consumption from the previous 23 hours up to the current time step (from $t-23$ to $t$), which serves as the input sequence for the LSTM neural network
For each of these 24 hourly consumption values, three contextual features were included:
  1. $f2)$
    the day of the year (ranging from 1 to 365)
  2. $f3)$
    the day of the week (ranging from 0 to 6)
  3. $f4)$
    the hour of the day (ranging from 0 to 23)
$f2,f3,f4$ (72 features in total) are fed into the dense neural network.
The architecture used, extensively studied in [19], is summarized in Fig. 3.
Figure 3: Hybrid LSTM neural network architecture, where nn labels the number of neurons.
Since our output is limited to a single hour of prediction, to predict multiple hours, we use the neural network recursively, where the output of each prediction becomes the input for the next.
The hybrid LSTM neural network was implemented using Tensorflow and Keras. The model was trained with a batch size of 8 and for 75 epochs. The model employed the Adam optimizer. The other settings are shown in Fig. 3. As demonstrated in [19], this architecture shows accurate performances for electrical load prediction.
The proposed prediction model works during the time window in which the EV is parked. Based on the household load predictions, the optimization problem with the objective function in (1) is solved to determine the optimal energy profiles. However, as time progresses, real consumption data become available for the earliest predicted hour. This allows for online energy profile corrections in case the prediction deviates from the actual household load. The main steps of this adaptive optimization are given in Algorithm 1. $t_{arr}$ and $t_{pick}$ are the EV arrival time and the pickup time, respectively. $HL_{p}$ is the predicted household load.
Algorithm 1 Energy optimization during parking mode
1:Start the parking mode ($t=t_{arr}$)
2:while $t<t_{pick}$ do
3:     if $t=t_{arr}$ OR $HL(t)\neq HL_{p}(t)$ then
4:         PREDICT $HL$ using the hybrid LSTM struc-
5:          tured in Fig. 3 (from $t+1$ to $t_{pick}-1$)
6:         OPTIMIZE $E^{V2G}$, $E^{V2H}$, $E^{G2H}$, and $E_{t}^{G2V}$
7:         (from $t$ to $t_{pick}-1$)
8:     $t=t+1$
9:End the parking mode ($t=t_{pick}$)
Note that in Algorithm 1, the condition of $HL(t)\neq HL_{p}(t)$ is used to adjust the energy flow based on household load predictions. Specifically, if the predicted household load deviates from the actual consumption, both the prediction (Step 4) and optimization (Step 5) are updated accordingly. In an ideal scenario with a perfect load predictor, this adjustment would not be necessary, and the optimization would be performed only once within the algorithm for each parking session.
II-D Data Sources and Simulation Data Generation
For the household load, the dataset from [20] has been used in this work. It contains household load data for single apartments in the US, each spanning one year. Five datasets from the state of Washington have been selected because they exhibit similar consumption patterns (shown in Fig. 4, where the datasets are concatenated to form a five-year dataset).
Figure 4: Concatenated household load for five apartments in the State of Washington, US. The red dashed vertical lines represent the transition between datasets.
Four of the five datasets were used for training, while one was used for testing, meaning 20% of the data was reserved for the model simulation. This last year of data, dedicated for the simulations, has an average hourly consumption of 0.9 kWh, resulting in a daily average consumption of 21.6 kWh.
As for the battery, the replacement cost $C_{rep}$ is defined as 111.5 €/kWh, the residual value is $C_{rv}=30\%C_{rep}$, the yearly discount rate is $i_{r}=10\%$ and the nominal battery life is $L=10$ years.
The daily EV usage and driving distances are modeled using a truncated Gaussian distribution. The EV is assumed to be picked up by the user between 6:00 and 10:00 am, with a mean pickup time of 8:00 am. The travel durations range from 7 to 11 hours, with a mean of 9 hours. The daily driving distance is assumed to lie between 30 and 40 km, with a mean of 35 km. We further assume that the driving distance is distributed linearly over the duration of the travel session.
The electricity price for the year 2022 for Sweden (SE3 area) is taken from ENTSO-E [21]. However, this represents the initial price ($p^{ini}$), i.e., the price without taxes. The final price $p$ used in the simulations has been calculated using
$p_{t}=p^{ini}_{t}\cdot(1+0.25)+0.006.$ (22)
A 25% markup is applied to the initial price to approximately account for Swedish value-added tax (VAT), and a fixed cost of 0.006 €/kWh is added to represent the energy tax.
To match the resolution of the energy price and household load datasets, simulations are conducted over one year with an hourly time step. The simulation started with the battery with an age of 60 days and the EV’s SoC at 60%. The complete control algorithm was developed in Python, with CasADi [22] serving as the optimization framework and the IPOPT solver applied for solving the nonlinear optimization problems.
III RESULTS AND DISCUSSION
Simulations were conducted for two distinct scenarios to evaluate user costs in the vehicle-home-grid integration:
  1. A.
    Vehicle-home-grid integration: With objective function in (1), this scenario aims to minimize the user energy costs and battery degradation by controlling the bidirectional energy flows among the vehicle, household, and electric grid.
  2. B.
    Unidirectional smart charging (benchmark): This scenario seeks to minimize costs, including battery degradation, without employing V2G and V2H technologies. Therefore, the objective function in (1) is replaced by
    $\min\sum_{t}(E^{G2V}_{t}+E^{G2H}_{t})\cdot p_{t}+BC_{t}+s_{t}.$ (23)
III-A Cost and Battery Degradation Analysis
With the price ratio $\gamma=1$, meaning the price for purchasing and selling energy is identical, Table I summarizes the user costs for the two scenarios. Here, the user’s final cost ($FC$) is calculated as the sum of the energy cost ($EC$) and battery degradation cost ($BC$). $BD^{cyc}$ is the sum of cycle aging under high temperature, low temperature, and high temperature with low SoC. $E_{batt}$ represents the total energy that flows in and out of the EV battery.
TABLE I: User costs and battery degradation for scenarios A and B.
$FC$ $EC$ $BC$ $BD$ $BD^{cal}$ $BD^{cyc}$ $E_{batt}$
[€] [€] [€] [%] [%] [%] [kWh]
A -1070.21 -1739.38 669.17 5.42 2.26 3.16 45030
B 1976.60 1549.12 427.48 3.46 2.72 0.75 4054
It can be observed that scenario A yields favorable performance: the negative energy cost ($EC$) indicates profit generated by selling electricity back to the grid via V2G. Correspondingly, the EV battery has degraded 5.42%. Overall, scenario A achieve an annual user profit ($FC$) of €1070.21.
In contrast, scenario B (benchmark) shows significantly lower battery energy flow ($E_{batt}$), limited to EV charging (G2V) and driving. Without energy sales (no V2G or V2H), scenario B results in significantly higher energy costs and minimal battery degradation due to both optimized smart charging and reduced battery usage.
By comparing scenarios A and B, it can be concluded that the vehicle-home-grid integration provides a substantial economic advantage of €3046.81 annually for the user. Specifically, scenario A degrades the battery by 1.96% more but reduces the energy costs by €3288.50 compared to scenario B.
III-B Sensitivity Analysis of Price Ratio $\gamma$
The results above assume a price ratio $\gamma=1$, commonly adopted as the most optimistic case. Namely, there is no price difference between the purchase and sale of energy. Different values of $\gamma$ may significantly affect the final cost $FC$ due to changes in the energy cost $EC$. Fig. 5 shows the final cost $FC$ for scenarios A and B across a range of $\gamma$ values from 0 to 1.
Figure 5: $FC$ for scenario A and B varying the price ratio $\gamma$.
As expected, in scenario B, the curve remains constant due to unidirectional charging, which does not involve V2G and thus is independent of the price ratio $\gamma$. In scenario A, for $\gamma=1$, $FC$ corresponds to the value shown in Table I. As $\gamma$ decreases, $FC$ increases because the reduced price difference between buying and selling makes V2G less profitable. For $\gamma=0.75$, $FC=0$. This indicates that using the EV for vehicle-home-grid integration eliminates EV-related costs, corresponding to an economic gain of €1976.60 that the user would have spent on unidirectional charging. However, for $\gamma<0.75$, $FC$ becomes positive, meaning that while the economic gain persists, no further profit is generated. When $\gamma=0$, V2G is no longer performed as it offers no benefit; however, V2H continues, contributing to self-consumption. This results in an economic gain (corresponding to a saving) of €425.48 compared to scenario B.
Overall, these results demonstrate that vehicle-home-grid integration consistently offers benefits: even in the worst-case scenario with $\gamma=0$, the user saves around €425 annually through bidirectional charging with their EV.
III-C Impact of Battery Capacity & Household Load Variations
Additional simulations are conducted to examine the influence of battery capacity ($E_{b}$ = [41, 61.5, 82, 102.5] kWh) and household loads ($\bm{HL}$, $4HL$, $8HL$). Here, $HL$ is the consumption of a standard apartment, as shown in Fig. 4, while $4HL$ and $8HL$ correspond to larger households (86.4 and 172.8 kWh/day, respectively). Fig. 6 illustrates the economic gain, defined as the difference in $FC$ between scenario B and A, for varying $E_{b}$, household loads, and $\gamma$.
Figure 6: Economic gain for varying battery capacities ($E_{b}$), household loads, and price ratios ($\gamma$).
When the consumption is $HL$ (solid lines), the curves for different $E_{b}$ values exhibit a clear increasing slope with $\gamma$, due to more effective V2G utilization. At low $\gamma$ values (less than 0.2), the economic gain decreases with larger batteries because any of considered battery capacities is sufficient to cover this $HL$, whereas larger batteries lead to increased battery replace costs, i.e., $C_{rep}$ in (4). Conversely, at higher $\gamma$ values, larger batteries result in greater economic gains due to enhanced V2G opportunities.
With increased household consumption ($4HL$ and $8HL$), the economic gain curves become flatter as $\gamma$ increases, indicating reduced V2G opportunities due to the higher household energy demand. Specifically, for $4HL$ (dashed lines), a 41 kWh battery is clearly limited and yields the smallest economic gain among all tested battery capacities. Moreover, in the $4HL$ case, smaller batteries offer slightly higher gains when $\gamma$ is low, while the opposite trend is observed at higher $\gamma$, consistent with the behavior seen in the $HL$ case. In the case of $8HL$ (dash-dot lines), the very high household load severely limits the availability of energy for grid export, effectively suppressing V2G utilization. Unlike the $HL$ and $4HL$ cases, a larger battery consistently yields higher economic gain across the entire $\gamma$ range. This is because larger batteries provide the capacity needed to handle the high household demand more flexibly, even if V2G is minimally utilized.
Additionally, for $\gamma=1$ and the same $E_{b}$, the economic gain remains constant across different household loads. This occurs because, with $\gamma=1$, there is no financial penalty on energy sales, and the user’s benefit depends solely on battery capacity. As a result, while a higher household load increases $FC$ in scenarios A and B, their economic difference remains unchanged for the same battery capacity.
Overall, results in Fig. 6 confirm that vehicle-home-grid integration consistently provides economic benefits across all tested cases. Specifically, regardless of the price ratios, battery capacities, and household loads, even in the least favorable conditions, the user achieves financial advantages compared to the unidirectional charging in scenario B.
IV CONCLUSION
This paper proposed a practical, real-time implementable optimization algorithm to manage the energy flows of vehicle-home-grid integration. The contributions first arise from the explicit consideration of nonlinear battery cycle and calendar aging, uncertainties in future household energy consumption, and a set of constraints in vehicle usage and battery dynamics. By utilizing a hybrid LSTM neural network to predict household loads and a detailed battery model to describe battery behavior, this algorithm results in minimized user energy and battery costs. This responds to a maximum annual economic gain of €3046.81 for a single user over the benchmark scenario with unidirectional smart charging.
The second major contribution is the conducted systematic analysis of various price ratios, battery capacities, and household loads within the optimization problem. We have found that
  • Reduced price ratios make V2G less profitable. However, even without performing V2G, V2H contributes to self-consumption and gives an annual economic gain of €425.48.
  • For standard apartments and small to moderate-sized houses, larger batteries reduce the economic gains at low price ratios but provide greater benefits at high price ratios.
  • For large household loads, larger batteries always lead to higher economic gains, as more efficient household energy management is allowed.